# Hierarchical Clustering With Confidence

### v1

@@ 3.1 Null hypothesis and conditional Type I error
For the observed realization of the data matrix $X=X_{o}$, suppose we observe
$\left\{M^{(t)}=M_{o}^{(t)}\right\},\quad\text{where}\quad M_{o}^{(t)}=\left(C^{(t)}_{1,o},C^{(t)}_{2,o}\right)$,
and for $j\in\{1,2\}$, $C^{(t)}_{j,o}$ is the observed value of ${C}^{(t)}_{j}$. We then define our null hypothesis as
@@
where under the null $H_{0}^{(t)}$, the data points belonging to the merged pair of clusters $C^{(t)}_{1,o}$ and $C^{(t)}_{2,o}$ are assumed to share the same mean parameter. In what follows, we derive an exact, closed-form, and valid p-value for testing this null hypothesis, treating the variance parameter $\sigma^{2}$ in (3) as unknown.
Before presenting our testing procedure, we first present the conditional Type I error guarantee associated with the p-values we construct for each merge. To formalize this guarantee, let us denote the sequence of random merges up to time $t$ as $\overline{M}^{(t)}=\left\{M^{(1)},M^{(2)},\dots,M^{(t)}\right\}$. Consistent with our notations, suppose we observe $\left\{\overline{M}^{(t)}=\overline{M}_{o}^{(t)}\right\}$ on our specific realization of the data matrix $X_{o}$, where
$\overline{M}_{o}^{(t)}=\left\{M_{o}^{(1)},M_{o}^{(2)},\ldots,M_{o}^{(t)}\right\},$
and $M_{o}^{(t)}$ is the observed value of $M^{(t)}$.
@@
As we show later, in constructing our $\alpha$-spending testing procedure, we rely on the conditional Type I error rate in (5), allowing for an adaptive choice of test level $\alpha=\alpha^{(t)}(\overline{M}^{(t)})$ at step $t$ that depends on all previous merges up to that step.
@@ 3.2 Test statistic and preliminaries
Given these definitions, it follows that: (i) $\nu_{M_{o}^{(t)}}^{T}X=\bar{X}_{C^{(t)}_{1,o}}-\bar{X}_{C^{(t)}_{2,o}}$ captures the differences between the means of clusters merged at step $t$; (ii) $B^{(t)}_{o}X$ is the projection of $X$ onto a rank-one subspace capturing the differences between the cluster means of the merged pair; (iii) $W^{(t)}_{o}X$ is the projection of $X$ onto a subspace with rank $N_{o}^{(t)}-2$ capturing the differences among observations within the two merged clusters; and (iv) the subspace that $W_{o}^{(t)}$ projects onto is orthogonal to that of $B_{o}^{(t)}$.
@@
Yun and Barber (2023) derive the distribution of this statistic under the null hypothesis (4) by conditioning on the clustering output, and show that it follows a truncated F distribution in the absence of additional randomization in the algorithm. In general, however, this null distribution is not available in closed form, as the truncation region typically lacks an explicit characterization, except in a few special cases. Consequently, Yun and Barber (2023) employ Monte Carlo approximations to compute their p-values, which cannot be calculated exactly. As emphasized earlier in the paper, we take a different approach to this problem by utilizing our randomization scheme to altogether bypass the need to describe the underlying conditioning event.
@@
The distribution of $R^{(t)}$ conditional on all merges up to step $t$ is difficult to characterize, since the merges made by a clustering algorithm depend on the entire data matrix rather than solely on the test statistic of interest. To this end, we state a few preliminaries, adopted from (Yun and Barber, 2023).
@@
$X(R^{(t)}_{o};\mathcal{A}^{(t)}_{o})=X_{o},$
where $R^{(t)}_{o}$ is as defined in (7).
The proof for this result can be found in Appendix B.3.
The next lemma shows that the test statistic $R^{(t)}$ given the auxiliary statistics has an $F$ distribution under $H_{0}^{(t)}$, provided that the clusters $(C^{(t)}_{1,o}$, $C^{(t)}_{2,o})$ are predefined, i.e., their dependence on data is ignored. Our main result, presented in the next section, further conditions this distribution on the sequence of greedy merges made by our randomized clustering algorithm to obtain the correct distribution of the test statistic.
Lemma 3.
@@
A detailed proof for this Lemma is deferred to Appendix B.4.
@@ 3.3 Constructing a conditionally valid p-value
Theorem 1 derives this conditional distribution, accounting for the highly data-dependent nature of the clustering process by conditioning on $\left\{\overline{M}^{(t)}=\overline{M}_{o}^{(t)}\right\}$, the sequence of greedy merges up to step $t$. The additional conditioning on the auxiliary statistics reduces the calculation of this conditional distribution to computing one-dimensional integrals, eliminating the need to marginalize over them in the density function.
@@
$\mathbb{F}^{(t)}(r;\mathcal{A}^{(t)}_{o},\overline{M}_{o}^{(t)})=\frac{\int_{0}^{r}\ell_{F_{p,(N_{o}^{(t)}-2)p}}(r^{\prime})\times\prod_{s=1}^{t}p^{(s)}\left(M_{o}^{(s)};X(r^{\prime};\mathcal{A}_{o}^{(t)})\right)dr^{\prime}}{\int_{0}^{\infty}\ell_{F_{p,(N_{o}^{(t)}-2)p}}(r^{\prime})\times\prod_{s=1}^{t}p^{(s)}\left(M_{o}^{(s)};X(r^{\prime};\mathcal{A}_{o}^{(t)})\right)dr^{\prime}},$
@@
$p^{(s)}\left(M_{o}^{(s)};X(r^{\prime};\mathcal{A}_{o}^{(t)})\right)=\mathbb{P}\left(M^{(s)}=M_{o}^{(s)}|X=X(r^{\prime};\mathcal{A}_{o}^{(t)})\right)$
is the sampling probability that $M^{(s)}=M_{o}^{(s)}$ given $X=X(r^{\prime};\mathcal{A}_{o}^{(t)})$, for $s\in\{1,2,\ldots,t\}$, as defined in (1).
Proof.
Let the conditional density of
$R^{(t)}\Big\lvert\mathcal{A}^{(t)}=\mathcal{A}_{o}^{(t)}$
at $r$ be denoted by $\ell_{R^{(t)}}(r\;\lvert\;\mathcal{A}^{(t)}=\mathcal{A}_{o}^{(t)}).$ Then, using Bayes’ rule, we observe that the density of the conditional distribution $R^{(t)}\Big\lvert\{\overline{M}^{(t)}=\overline{M}_{o}^{(t)},\mathcal{A}^{(t)}=\mathcal{A}_{o}^{(t)}\},$ evaluated at $r$, is proportional to
Using Lemma 3, under the null $H_{0}^{(t)}$, it follows that
Furthermore, since the merges were sampled according to the probabilities in (1), $\{R^{(t)}=r,\mathcal{A}^{(t)}=\mathcal{A}_{o}^{(t)}\}=\{X=X(r;\mathcal{A}_{o}^{(t)})\}$ by the definition in Lemma 2, and the sampling processes were independent given the data, we have that
$=\prod_{s=1}^{t}\mathbb{P}\left(M^{(s)}=M_{o}^{(s)}|X=X(r;\mathcal{A}_{o}^{(t)})\right)=\prod_{s=1}^{t}p^{(s)}\left(M_{o}^{(s)};X(r;\mathcal{A}_{o}^{(t)})\right).$
Combining (13) with (14), we obtain that the expression in (12) takes the value $\ell_{F_{p,(N_{o}^{(t)}-2)p}}(r)\times\prod_{s=1}^{t}p^{(s)}\left(M_{o}^{(s)};X(r;\mathcal{A}_{o}^{(t)})\right).$ That is, the conditional density of interest is given by
$\frac{\ell_{F_{p,(N_{o}^{(t)}-2)p}}(r)\times\prod_{s=1}^{t}p^{(s)}\left(M_{o}^{(s)};X(r;\mathcal{A}_{o}^{(t)})\right)}{\int_{0}^{\infty}\ell_{F_{p,(N_{o}^{(t)}-2)p}}(r^{\prime})\times\prod_{s=1}^{t}p^{(s)}\left(M_{o}^{(s)};X(r^{\prime};\mathcal{A}_{o}^{(t)})\right)dr^{\prime}}.$
Next, we present the conditional p-value, denoted by ${\textbf{P}}^{(t)}={\textbf{P}}^{(t)}\left(R^{(t)};\mathcal{A}^{(t)}\right)$, which is a function of the test statistic and the auxiliary statistics. Let
${\textbf{P}}^{(t)}_{o}={\textbf{P}}^{(t)}\left(R^{(t)}_{o};\mathcal{A}^{(t)}_{o}\right)$
denote its value computed on the observed data $X_{o}$, defined as
${\textbf{P}}^{(t)}_{o}=\mathbb{P}_{H_{0}^{(t)}}\left(R^{(t)}\geq R^{(t)}_{o}\;\Big\lvert\;\left\{\overline{M}^{(t)}=\overline{M}_{o}^{(t)},\mathcal{A}^{(t)}=\mathcal{A}_{o}^{(t)}\right\}\right),$ (15)
@@
A proof for this result is included in Appendix B.5. Simulations in Section 5.2 demonstrate that the p-values constructed in this section not only control the Type I error rate at the nominal level, but also exhibit higher power than existing selective inferential methods for producing p-values at a fixed dendrogram level.
We conclude this section with several remarks on the test developed in this section. As shown in Corollary 1, a conditionally valid p-value is derived from the CDF of the conditional distribution in Theorem 1. Unlike earlier non-randomized selective inference methods, however, this CDF (see the proof of Theorem 1) does not require an explicit linkage- and dissimilarity-specific characterization of the conditioning event. Instead, the correction for double dipping is obtained directly from the sampling probabilities of the observed sequence of merges, defined in (1), which already encode the linkage dissimilarity. Consequently, computing the CDF, and thus the p-value, requires no sampling or resampling based approximations. For example, Yun and Barber (2023) uses Monte Carlo approximations to compute p-values when more than two clusters are present, i.e., at any intermediate step of the merging process. Similarly, Gao et al. (2024) rely on Monte Carlo approximations to compute p-values for complete-linkage clustering, and do not compute p-values for linkages such as the minimax linkage or for dissimilarities outside the class of squared Euclidean distances. By contrast, our method avoids such approximations and computes p-values directly from the sampling probabilities of our randomization scheme.
Finally, in Appendix A.1, we also develop conditional p-values for the type of test proposed by Gao et al. (2024) assuming a normal model with a known non-spherical covariance matrix $\Sigma$. By using a slightly different decomposition of the data matrix into the test statistic and appropriately defined auxiliary statistics, the CDF of the relevant conditional distribution is derived from the sampling probabilities in our randomization scheme, analogous to the result in Theorem 1. This CDF yields valid p-values with the conditional Type I error guarantee in (5).
@@ 4 Using conditional p-values to choose the number of clusters
We develop an adaptive alpha-spending procedure Foster and Stine (2008); Aharoni and Rosset (2014) that leverages the conditionally valid p-values from the previous section to estimate the number of clusters. Specifically, we begin with a pre-defined sequence $\{\alpha_{1},\dots,\alpha_{n-1}\}$ satisfying $\sum_{j=1}^{n-1}\alpha_{j}=\alpha$, where (n1)(n-1) is the total number of merges. At each merge in the dendrogram, we apply the test described in Section 3 and compute the corresponding p-value ${\textbf{P}}^{(t)}_{o}$ for the null hypothesis specified in (4). We then compare this p-value to $\alpha^{(t)}$, which is selected from $\mathcal{S}$, a subset of unused elements of $\{\alpha_{1},\dots,\alpha_{n-1}\}$ at step $t$, and terminate the sequential testing procedure as soon as there is sufficient evidence against a merge made by Algorithm 1.
The choice of $\alpha^{(t)}$ in our procedure is allowed to depend adaptively on the sizes of the clusters being merged, and, by design, also on the $\alpha$-budget already allocated to the previous merges up to time $t$. In particular, we design a prudent, adaptive allocation of $\alpha$-values that enables the use of more liberal thresholds for merges involving larger clusters, leading to a more powerful procedure.
The full procedure is summarized in Algorithm 2. In our implementation, we introduce a minimum cluster size cutoff $n_{\text{min}}$ such that if any cluster has size below this threshold, we do not perform a test and the corresponding merges are automatically accepted. This includes, trivially, the case in which the clusters are singletons, for which the naive F test, and consequently the appropriately adjusted test in Section 3 does not apply. Additionally, if the minimum of the two cluster sizes is larger than $n_{\text{min}}$ but smaller than $n^{*}$ at any step $t$, we assign a smaller $\alpha^{(t)}$ from the list of unused elements of $\{\alpha_{1},\dots,\alpha_{n-1}\}$ at step $t$. If $n^{*}$ is chosen to be less than or equal to $n_{\text{min}}$, this step is skipped. This type of adaptive allocation of $\alpha$-values is consistent with our goal of using more liberal thresholds for merges involving larger clusters and therefore obtain a more powerful procedure.
We next develop a probabilistic guarantee for our algorithm against overestimation. To introduce this result, suppose $K^{*}$ is the true number of clusters in the data, and let $t^{*}=n-K^{*}+1$. Denote by $\mathcal{T}$ the event that the true cluster structure is preserved when the randomized hierarchical clustering algorithm is executed up to step $t^{*}$, i.e., that no cross-cluster merges occur before $t^{*}$.
@@
Suppose Algorithm 2 is executed with total significance level $\alpha$ and a pre-specified sequence $\alpha_{1},\alpha_{2},\dots,\alpha_{n-1}$} satisfying $\sum^{n-1}_{j=1}\alpha_{j}=\alpha$. Then the $\alpha$-spending procedure summarized in Algorithm 2 guarantees
$\mathbb{P}\left(\{\widehat{K}>K^{*}\}\cap\mathcal{T}\right)\leq\alpha.$
A proof for this result is included in Appendix B.6. It proceeds by demonstrating that this type of sequential testing procedure satisfies family-wise error rate (FWER) control, which is shown in Lemma 6.
@@
1:Input: Significance level $\alpha$, pre-defined sequence $\{\alpha_{1},\alpha_{2},\dots,\alpha_{n-1}\}$ such that $\sum_{j=1}^{n-1}\alpha_{j}=\alpha$, merge sequence $\overline{M}_{o}^{(t)}=\{(C^{(t)}_{1,o},C^{(t)}_{2,o})\}_{t=1}^{n-1}$ obtained from Algorithm 1 with randomization level $\tau$, testable cluster size cutoff $n_{\text{min}}$, alpha-spending threshold cluster size $n^{*}$
2:Initialize: Available alpha set $\mathcal{S}=\{\alpha_{1},\alpha_{2},\dots,\alpha_{n-1}\}$
@@
4:  Conduct the greedy merge $M_{o}^{(t)}=(C^{(t)}_{1,o},C^{(t)}_{2,o})$
5:  if $\min(n_{1},n_{2})\leq n_{\text{min}}$ then
6:   No test is performed, and we proceed to the next merge
7:  else
8:   if $\min(|C^{(t)}_{1,o}|,|C^{(t)}_{2,o}|)\leq n^{*}$ then
9:     Select test level $\alpha^{(t)}\leftarrow\min(\mathcal{S})$ $\triangleright$ More conservative for small clusters
10:   else
11:     Select test level $\alpha^{(t)}\leftarrow\max(\mathcal{S})$ $\triangleright$ More power for larger clusters
12:   end if
13:   Conduct hypothesis test (4) by constructing the p-value ${\textbf{P}}^{(t)}_{o}$ in (15)
14:  end if
15:  Update available alpha set: $\mathcal{S}\leftarrow\mathcal{S}\setminus\{\alpha^{(t)}\}$
16:  if ${\textbf{P}}^{(t)}_{o}<\alpha^{(t)}$ then
17:   Set estimated number of clusters $\widehat{K}=n-t+1$
18:   Output: $\widehat{K}$
19:   Break
20:  end if
21:end for
22:If no rejection: Set $\widehat{K}=1$
23:Output: $\widehat{K}$
The numerical experiments on simulated and real data in Sections 5.3 and 5.4 investigate the potential of the proposed $\alpha$-spending procedure for estimating the dendrogram cut height. The resulting estimate of the number of clusters can then be used to assess standard metrics for clustering stability across repeated runs of our randomized algorithm or clustering algorithms with other forms of randomization. Across a variety of simulation settings, we find that our testing procedure for estimating the number of clusters yields a powerful clustering method that outperforms standard (i.e., non-randomized) clustering approaches paired with the popular gap statistic (Tibshirani et al., 2001).

### v2

@@ 3.1 Null hypothesis and conditional Type I error
A particular realization of Algorithm 1 on $X_{o}$ (an observed realization of the random matrix $X$), yields a fixed sequence of merges up to time $t$, which we will denote $\overline{M}_{o}^{(t)}=\left(M_{o}^{(1)},M_{o}^{(2)},\ldots,M_{o}^{(t)}\right).$ At step $t$, the chosen merge $M_{o}^{(t)}=\left(C^{(t)}_{1,o},C^{(t)}_{2,o}\right)$ gives rise to the null hypothesis
@@
which states that the data points belonging to the merged pair of clusters $C^{(t)}_{1,o}$ and $C^{(t)}_{2,o}$ share the same mean parameter.
In what follows, we derive an exact, closed-form, and valid p-value for testing this null hypothesis, treating the variance parameter $\sigma^{2}$ in (5) as unknown. Before presenting our testing procedure, we first present the conditional Type I error guarantee associated with the p-values we construct for each merge.
@@
In words, ${\textbf{P}}^{(t)}$ follows a uniform distribution when we restrict attention to those runs of our randomized clustering algorithm on $X$ that produce the observed sequence of merges $\overline{M}_{o}^{(t)}$. As we show later, in constructing our $\alpha$-spending testing procedure, we rely on the conditional Type I error rate in (7), allowing for an adaptive choice of test level $\alpha=\alpha^{(t)}(\overline{M}^{(t-1)})$ at step $t$ that depends on all previous merges up to that step.
@@ 3.2 Test statistic and preliminaries
Given these definitions, it follows that: (i) $\nu_{M_{o}^{(t)}}^{T}X=\bar{X}_{C^{(t)}_{1,o}}-\bar{X}_{C^{(t)}_{2,o}}$ captures the differences between the means of clusters merged at step $t$; (ii) $B^{(t)}_{o}X$ is the projection of $X$ onto a rank-one subspace capturing the differences between the cluster means of the merged pair; (iii) $W^{(t)}_{o}X$ is the projection of $X$ onto a subspace with rank $N_{o}^{(t)}-2$ capturing the differences among observations within the two merged clusters relative to their cluster means; and (iv) the subspace that $W_{o}^{(t)}$ projects onto is orthogonal to that of $B_{o}^{(t)}$.
@@
For traditional (non-randomized) hierarchical clustering, Yun and Barber [2023] derive the distribution of this statistic under the null hypothesis (6) by conditioning on the clustering output, showing that it follows a truncated F distribution. In general, however, this null distribution is not available in closed form, as the truncation region typically lacks an explicit characterization, except in a few special cases. Consequently, Yun and Barber [2023] employ Monte Carlo approximations to compute their p-values, which cannot be calculated exactly. As emphasized earlier in the paper, we take a different approach to this problem by utilizing our randomization scheme to altogether bypass the need to describe the underlying conditioning event.
@@
The distribution of $R^{(t)}$ conditional on all merges up to step $t$ is difficult to characterize, since the merges made by a clustering algorithm depend on the entire data matrix rather than solely on the test statistic of interest. To this end, we state a few preliminaries, adopted from Yun and Barber [2023].
@@
[…]
The next lemma shows that the test statistic $R^{(t)}$ given the auxiliary statistics has an $F$ distribution under $H_{0}^{(t)}$, provided that the clusters $(C^{(t)}_{1,o}$, $C^{(t)}_{2,o})$ are predefined, i.e., their dependence on data is ignored.
Lemma 2 (Based on Yun and Barber [2023]).
@@
[…]
Our main result, presented in the next section, further conditions this distribution on the sequence of merges made by our randomized clustering algorithm to obtain the correct distribution of the test statistic and construct a p-value that satisfies (7).
@@ 3.3 Constructing a conditionally valid p-value
Theorem 1 derives this conditional distribution, accounting for the highly data-dependent nature of the clustering process by conditioning on $\left\{\overline{M}^{(t)}=\overline{M}_{o}^{(t)}\right\}$, the sequence of merges up to step $t$. The additional conditioning on the auxiliary statistics reduces the calculation of this conditional distribution to computing one-dimensional integrals, eliminating the need to marginalize over them in the density function.
@@
$\mathbb{F}^{(t)}(r;\mathcal{A}^{(t)}_{o},\overline{M}_{o}^{(t)})=\frac{\int_{0}^{r}\ell_{F_{p,(N_{o}^{(t)}-2)p}}(r^{\prime})\times\prod_{s=1}^{t}p^{(s)}\left(M_{o}^{(s)};\mathcal{M}_{o}^{(s)},X(r^{\prime};\mathcal{A}_{o}^{(t)})\right)dr^{\prime}}{\int_{0}^{\infty}\ell_{F_{p,(N_{o}^{(t)}-2)p}}(r^{\prime})\times\prod_{s=1}^{t}p^{(s)}\left(M_{o}^{(s)};\mathcal{M}_{o}^{(s)},X(r^{\prime};\mathcal{A}_{o}^{(t)})\right)dr^{\prime}},$
@@
$p^{(s)}\left(M_{o}^{(s)};\mathcal{M}_{o}^{(s)},X(r^{\prime};\mathcal{A}_{o}^{(t)})\right)=\mathbb{P}\left(M^{(s)}=M_{o}^{(s)}|\overline{M}^{(s-1)}=\overline{M}_{o}^{(s-1)},X=X(r^{\prime};\mathcal{A}_{o}^{(t)})\right)$
is the sampling probability that $M^{(s)}=M_{o}^{(s)}$ given $X=X(r^{\prime};\mathcal{A}_{o}^{(t)})$ and the merge history $\overline{M}^{(s-1)}=\overline{M}_{o}^{(s-1)}$, for $s\in\{1,2,\ldots,t\}$, as defined in (2).
[…]
Next, we present the conditional p-value, denoted by ${\textbf{P}}^{(t)}={\textbf{P}}^{(t)}\left(R^{(t)};\mathcal{A}^{(t)}\right)$, which is a function of the test statistic and the auxiliary statistics. Let ${\textbf{P}}^{(t)}_{o}={\textbf{P}}^{(t)}\left(R^{(t)}_{o};\mathcal{A}^{(t)}_{o}\right)$ denote its value computed on the observed data $X_{o}$, defined as
${\textbf{P}}^{(t)}_{o}=\mathbb{P}_{H_{0}^{(t)}}\left(R^{(t)}\geq R^{(t)}_{o}\;\Big\lvert\;\overline{M}^{(t)}=\overline{M}_{o}^{(t)},\mathcal{A}^{(t)}=\mathcal{A}_{o}^{(t)}\right),$ (14)
@@
[…]
3.4 Computational benefits of our randomization scheme
Here we highlight the benefits of our randomization scheme by contrasting the p-values in (14) with those derived from a truncated F distribution using the traditional clustering algorithm. Introducing some notations, let $\mathbb{T}^{(t)}(r;\mathcal{A}^{(t)}_{o},\overline{M}_{o}^{(t)})$ denote the CDF of an $F_{p,(N^{(t)}_{o}-2)p}$ random variable truncated to the same observed sequence of merges $\left\{\overline{M}^{(t)}=\overline{M}_{o}^{(t)}\right\}$, which is given by
$\mathbb{T}^{(t)}(r;\mathcal{A}^{(t)}_{o},\overline{M}_{o}^{(t)})=\frac{\int_{0}^{r}\ell_{F_{p,(N_{o}^{(t)}-2)p}}(r^{\prime})\times\prod_{s=1}^{t}\mathbbm{1}\{M_{o}^{(s)}\in\Psi^{*(s)}(r^{\prime})\}dr^{\prime}}{\int_{0}^{\infty}\ell_{F_{p,(N_{o}^{(t)}-2)p}}(r^{\prime})\times\prod_{s=1}^{t}\mathbbm{1}\{M_{o}^{(s)}\in\Psi^{*(s)}(r^{\prime})\}dr^{\prime}},$ (15)
where $\Psi^{*(s)}(r^{\prime})=\displaystyle\operatorname*{arg\,min}_{M\in\mathcal{M}_{o}^{(s)}}d(M;X(r^{\prime};\mathcal{A}^{(t)}_{o}))$. This yields the p-value ${\textbf{P}}^{(t)}_{\text{NR},o}=1-\mathbb{T}^{(t)}(R^{(t)}_{o};\mathcal{A}^{(t)}_{o},\overline{M}_{o}^{(t)})$, constructed without external randomization, which coincides with the p-value in Theorem 2 of Yun and Barber [2023] if the sequence of merges leading to the merge $M_{o}^{(t)}$ at step $t$ is unique.
Proposition 2 shows that, as the amount of randomization becomes smaller, i.e., as $\tau^{*}\rightarrow 0$, our p-values converge to ${\textbf{P}}^{(t)}_{\text{NR},o}=1-\mathbb{T}^{(t)}(R^{(t)}_{o};\mathcal{A}^{(t)}_{o},\overline{M}_{o}^{(t)})$.
Proposition 2.
It holds that, as $\tau^{*}\rightarrow 0$, ${\textbf{P}}^{(t)}_{o}(\tau^{*}){\rightarrow}{\textbf{P}}^{(t)}_{\text{NR},o}$.
[…]
The p-value ${\textbf{P}}^{(t)}_{\text{NR},o}$ is, however, not available in closed form, since the event $\left\{\overline{M}^{(t)}=\overline{M}_{o}^{(t)}\right\}$ (see the indicators in the numerator and denominator of (15)) lacks an analytical description except for certain special linkages and cases. For example, Yun and Barber [2023] uses Monte Carlo approximations to compute p-values when more than two clusters are present, i.e., at any intermediate step of the merging process. Similarly, Gao et al. [2024] rely on Monte Carlo approximations to compute p-values for complete-linkage clustering, and do not compute p-values for linkages such as the minimax linkage or for dissimilarities outside the class of squared Euclidean distances. Our p-values, on the other hand (see the CDF in Theorem 1), do not require an explicit, linkage- and dissimilarity-specific characterization of the conditioning event. Instead, the correction for double dipping is derived directly from the sampling probabilities of the observed sequence of merges, defined in (2), which already encode the linkage dissimilarity and are available in closed-form.
[…]
@@ 4 Using conditional p-values to choose the number of clusters
We develop an adaptive $\alpha$-spending procedure [Foster and Stine, 2008, Aharoni and Rosset, 2014] that leverages the conditionally valid p-values from the previous section to estimate the number of clusters. Specifically, we begin with a pre-defined sequence $\{\alpha_{1},\dots,\alpha_{n-1}\}$ satisfying $\sum_{j=1}^{n-1}\alpha_{j}=\alpha$, where (n1)(n-1) is the total number of merges. At each merge in the dendrogram, we compute the p-value ${\textbf{P}}^{(t)}_{o}$ for the null hypothesis specified in (6) and compare ${\textbf{P}}^{(t)}_{o}$ to a threshold $\alpha^{(t)}$ selected from $\mathcal{S}$, a subset of unused $\alpha$ values at step $t$. The procedure terminates as soon as there is sufficient evidence against a merge made by Algorithm 1.
The choice of $\alpha^{(t)}$ is allowed to depend adaptively on the sizes of the clusters being merged, and, by design, also on the $\alpha$-budget already allocated up to time $t$, which permits more liberal thresholds for merges involving larger clusters, leading to a more powerful procedure.
The full procedure is summarized in Algorithm 2. In our implementation, we introduce a minimum cluster size cutoff $n_{\text{min}}$ such that if any cluster has size below this threshold, the corresponding merges are automatically accepted without testing. Additionally, if the smaller cluster size lies between $n_{\text{min}}$ and the $\alpha$-spending threshold cluster size $n^{*}$ at any step $t$, we assign a smaller $\alpha^{(t)}$ from the list of unused elements of $\mathcal{S}$. If $n^{*}$ is chosen to be less than or equal to $n_{\text{min}}$, this step is skipped. This type of adaptive allocation of $\alpha$-values is consistent with our goal of using more liberal thresholds for merges involving larger clusters with the goal of obtaining a more powerful procedure.
We next develop a probabilistic guarantee for our algorithm against overestimation. Suppose $K^{*}$ is the true number of clusters in the data, and let $t^{*}=n-K^{*}+1$. Denote by $\mathcal{T}$ the event that the true cluster structure is preserved when the randomized hierarchical clustering algorithm is executed up to step $t^{*}$, i.e., that no cross-cluster merges occur before $t^{*}$.
@@
Suppose Algorithm 2 is executed with total significance level $\alpha$ and a pre-specified sequence $\alpha_{1},\alpha_{2},\dots,\alpha_{n-1}$} satisfying $\sum^{n-1}_{j=1}\alpha_{j}=\alpha$. Then the $\alpha$-spending procedure summarized in Algorithm 2 guarantees $\mathbb{P}\left(\{\widehat{K}>K^{*}\}\cap\mathcal{T}\right)\leq\alpha$.
[…]
The numerical experiments on simulated and real data in Sections 5.3 and 5.4 investigate the potential of the proposed $\alpha$-spending procedure for estimating the dendrogram cut height. The resulting estimate of the number of clusters can then be used to assess standard metrics for clustering stability across repeated runs of our randomized algorithm or clustering algorithms with other forms of randomization. Across a variety of simulation settings, we find that our testing procedure for estimating the number of clusters yields a powerful clustering method that outperforms standard (i.e., non-randomized) clustering approaches paired with the popular gap statistic [Tibshirani et al., 2001].
@@
1:Input: Significance level $\alpha$, pre-defined sequence $\{\alpha_{1},\alpha_{2},\dots,\alpha_{n-1}\}$ such that $\sum_{j=1}^{n-1}\alpha_{j}=\alpha$, merge sequence $\overline{M}_{o}^{(n-1)}=\{(C^{(t)}_{1,o},C^{(t)}_{2,o})\}_{t=1}^{n-1}$ obtained from Algorithm 1 with randomization level $\tau^{*}$, testable cluster size cutoff $n_{\text{min}}$, $\alpha$-spending threshold cluster size $n^{*}$
2:Initialize: Available set $\mathcal{S}=\{\alpha_{1},\alpha_{2},\dots,\alpha_{n-1}\}$
@@
4:  if $\min(|C^{(t)}_{1,o}|,|C^{(t)}_{2,o}|)\leq n_{\text{min}}$ then
5:   No test is performed, and we proceed to the next merge
6:  else
7:   if $\min(|C^{(t)}_{1,o}|,|C^{(t)}_{2,o}|)\leq n^{*}$ then
8:     Select test level $\alpha^{(t)}\leftarrow\min(\mathcal{S})$ $\triangleright$ More conservative for small clusters
9:   else
10:     Select test level $\alpha^{(t)}\leftarrow\max(\mathcal{S})$ $\triangleright$ More power for larger clusters
11:   end if
12:   Test the null hypothesis in (6) by constructing the p-value ${\textbf{P}}^{(t)}_{o}$ in (14)
13:  end if
14:  Update available alpha set: $\mathcal{S}\leftarrow\mathcal{S}\setminus\{\alpha^{(t)}\}$
15:  if ${\textbf{P}}^{(t)}_{o}<\alpha^{(t)}$ then
16:   Set estimated number of clusters $\widehat{K}=n-t+1$
17:   Output: $\widehat{K}$
18:   Break
19:  end if
20:end for
21:If no rejection: Set $\widehat{K}=1$
22:Output: $\widehat{K}$